OpenStudy (anonymous):

What is the last step when solving the equation below for x? 4x - 1 = 39

OpenStudy (anonymous):

4x = 39 + 1 x = 40/4

OpenStudy (anonymous):

x= 10 I believe

OpenStudy (gabebae):

4x-1=39 -1 -1 --------- 4x =38 - - 4 4 --------- x = 9.5 So the last step it dividing 38 by 4

OpenStudy (gabebae):

oops I subtracted 1

OpenStudy (gabebae):

My bad

OpenStudy (anonymous):

hehe

OpenStudy (anonymous):

@Lily2913 is right

OpenStudy (gabebae):

Ya, the answer is 10. I'm sorry about that

OpenStudy (anonymous):